OpenStudy (anonymous):

A right triangle has one side of length 11 and a hypotenuse of length 12. How long is the third side of the triangle? The square root of 265 1 23 The square root of 23

OpenStudy (anonymous):

Use Pythagorean Theorem: \[a^2 + b^2 = c^2\] Here a is one side , b is another side and c is hypotenuse..

OpenStudy (anonymous):

\[11^2 + b^2 = 12^2 \implies b^2 = 12^2 - 11^2 = ??\]

OpenStudy (anonymous):

@joyvise what is \(12^2\) ??

OpenStudy (anonymous):

144

OpenStudy (unklerhaukus):

...and \(11^2\)?

OpenStudy (anonymous):

Yes and what is \(11^2\) ??

OpenStudy (anonymous):

121

OpenStudy (anonymous):

Then what is \(144 - 121 = ??\)

OpenStudy (anonymous):

23

OpenStudy (anonymous):

\[b^2 = 23\] Now take the square root both the sides and find b..

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

Yeah but can you tell what is the answer ??

OpenStudy (anonymous):

i got 4.79

OpenStudy (anonymous):

But that is not in the choices ??

OpenStudy (anonymous):

the answer would just be square root of 23

OpenStudy (anonymous):

Say the answer according to the choices given to you..

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

ok